ADNAN MUHAMMED

Grade 12,

If lx+my+3+0 and 2y=3x-1 represent the same line then find value of l and m

If lx+my+3+0 and 2y=3x-1 represent the same line then find value of l and m

Grade:11

1 Answers

Arun
5 years ago
Dear Anika
 
If both lines are equal
Then
l/3 = m/-2 = -1/3
 
l = -1
m = 2/3
 
Hope it helps
 
 
Regards
Arun (askIITians forum expert)